# Enhancing Quantum Key Distribution with Entanglement Distillation and Classical Advantage Distillation

### v1

@@ IV.1.2 BB84 key rates
To consider the performance of the ED+AD scheme for the key distribution task, we calculate the asymptotic BB84 key rates [17, 23] for different protocols as follows.
@@
where $p_{\text{ED}}$ is the success probability for the entanglement distillation protocol, $p_{\text{AD}}$ the success probability for the advantage distillation protocol, $p_{\text{bit}}(p_{\text{phase}})$ the probability of the bit-flip (phase-flip) error for the distilled state, and $H(\cdot)$ is the usual binary entropy function. If the ED-only protocol is used, $p_{\text{AD}}$ and $n$ are set to 1.
@@
In the low noise regime, the ED-only protocols and $m$-$n$-$1$ protocols with smaller $m$ and $n$ perform better. This is due to the overhead incurred by the distillation. However, in the high noise regime, the advantage of ED+AD protocols begins to emerge. In the zoomed-in version of Fig. 5, if the input fidelity is lower than $0.68$, the ED-only protocols cannot achieve finite key rates, whereas several ED+AD protocols can still achieve finite key rates.
Rather surprisingly, in this noise regime, the protocol $2$-$3$-$1$ performs better than protocol $3$-$2$-$1$. This implies that in the case of depolarizing noise and using a quantum computer that can process at most $3$ qubits as input, using a smaller entanglement distillation ($m=2$) more often ($n=3$) is more advantageous than using a larger entanglement distillation protocol ($m=3$) and performing advantage distillation fewer times ($n=2$). The reason for the rather counter-intuitive situation is as follows: unlike the achievable fidelity calculation, when calculating the achievable key rates, the success probability of the distillation is also important. Although the $3$-$2$-$1$ protocols can, in principle, distill a more entangled state than $2$-$3$-$1$ protocols, the lower overall success probability makes the former less advantageous for performing QKD. In fact, this poses good news for experimental implementation since the larger entanglement distillation protocols require more qubits and longer coherence time in the quantum computer.
@@ IV.2.2 BB84 key rates
In terms of achievable key rates, in all noise regimes, the ED-only protocol $2$-$1$ outperforms all other more complicated protocols (Fig. 7). The special structure of the dephasing noise eliminates the need for complicated distillation protocols such as the $3$-$3$-$1$. While these protocols can produce high-quality entanglement (Fig. 6), the low success probability makes them less efficient compared to ED-only protocols. The results here demonstrate that depending on the nature of the noise in the input state as well as the strength of the noise, very different distillation protocols should be used in order to achieve optimal performance of entanglement-based QKD.
@@ V Conclusion
We presented a scheme for improving entanglement-based quantum key distribution, based on entanglement distillation and advantage distillation. We find both increased secret-key rates and increased resilience against noise, while not imposing any additional requirements on the quantum resources. In particular, our protocol has the potential to achieve a non-zero key rate even at a noise regime where using entanglement distillation only leads to a vanishing key rate for the case of depolarizing noise.
Our approach also works for more general Pauli noise models — if one possesses knowledge about the initial noisy state shared by two parties, our framework can be immediately utilized to find the optimal ED+AD protocol for performing entanglement-based QKD. Applying our tools to the case of pure-dephasing noise, we found that the basic $2$-$1$ protocol achieves better key rates than more sophisticated protocols, further highlighting the practical implications of our results on near-term quantum communication.
Our work suggests several possible future directions of study. First, due to the still rather large number of all possible ED+AD protocols, studying the $m$-$n$-$1$ protocol for large $m$ and $n$ is still formidable. A discrete optimization algorithm, rather than enumerating all possible protocols, should be employed to study larger protocols. Second, we have not considered the case of more general $m$-$n$-$o$ ED+AD protocols ($n>o>1$). In this more general case, one would need to consider all possible $n$-$o$ advantage distillation protocols to obtain optimal protocols. Finally, in practice one has to deal with non-identical input states and the effect of imperfect quantum operations during entanglement distillation [14]. We believe the framework built here can be extended to take the above considerations in mind, which in turn will yield more efficient entanglement-based QKD protocols and applications in the near future.
@@ Appendix A Binary picture and the symplectic representation of Clifford operators
For enumerating the bi-local Clifford protocols, it will be convenient to define an equivalence relation on such protocols; equivalent protocols have equivalent distillation statistics, meaning the success probability and output state (up to local rotations) are the same. Here we review the required mathematical framework. The material here follows closely Ref. [13].
In the study of Clifford distillation protocols, the following binary vector representation for the Pauli group will be useful. The convention here follows Ref. [25] and Ref. [13]. We define the following mapping between binary vectors and single-qubit Pauli operators as follows,
@@
Nevertheless, when one considers using Clifford gates to perform entanglement distillation, not all choices of Clifford operators lead to different distillation statistics. Two protocols have the same distillation statistics if the output fidelity $F=\langle\Phi^{+}|\rho_{distilled}|\Phi^{+}\rangle$ and the success probability are the same, and other components with respect to the Bell basis ($p_{X}=\langle\Psi^{+}|\rho_{\text{distilled}}|\Psi^{+}\rangle,p_{Y}=\langle\Psi^{-}|\rho_{\text{distilled}}|\Psi^{-}\rangle,p_{Z}=\langle\Phi^{-}|\rho_{\text{distilled}}|\Phi^{-}\rangle$) are the same up to some permutation. In Ref. [13], the equivalence classes (up to the same distillation statistics) of the symplectic group are characterized, which has a much smaller size than the original symplectic group. Let ${D}_{n}$ be an element in the symplectic group $\text{Sp}(2n,\mathbb{Z}_{2})$ which, when left multiplied to any Clifford operator $C_{n}$ in the distillation protocol, does not change the distillation statistics. Such ${D}_{n}$ form a subgroup $\mathcal{D}_{n}$. When one wants to enumerate all possible bi-local Clifford protocols for a given input state, it suffices to enumerate one element in each cosets of $\mathcal{D}_{n}$ in $\text{Sp}(2n,\mathbb{Z}_{2})$, which has size equal to the associated subgroup index
@@
Appendix B Comparison with DEJMPS-based protocols
@@
Figure 8: Illustration of a $3$-$1$ DEJMPS protocol. The $3$-$1$ DEJMPS protocol is the recursive version of standard DEJMPS protocol. The first unmeasured pair is the output of the protocol.
@@
Appendix C Repetition code based distillation on Werner states
@@ Acknowledgements.
We thank Hiroki Takahashi for stimulating discussions. This project was supported by the JST Moonshot R&D program under Grant JPMJMS226C.

### v2

@@ IV.1.2 BB84 key rates
To consider the performance of the ED+AD scheme for key distribution, we calculate the asymptotic BB84 key rates [20, 26] for different protocols as follows.
@@
where $p_{\text{ED}}$ ($p_{\text{AD}}$) is the success probability for the entanglement distillation protocol (the success probability for the advantage distillation protocol), $p_{\text{bit}}(p_{\text{phase}})$ the probability of the bit-flip (phase-flip) error for the distilled state, and $H(\cdot)$ is the usual binary entropy function. If the ED-only protocol is used, $p_{\text{AD}}$ and $n$ are set to 1.
@@
In the low noise regime, the ED-only protocols and $m$-$n$-$1$ protocols with smaller $m$ and $n$ perform better. This is due to the overhead incurred by the distillation. However, in the high noise regime, the advantage of ED+AD protocols begins to emerge. In the zoomed-in version of Fig. 5, if the input fidelity is lower than around $0.67$, the ED-only protocols cannot achieve finite key rates, whereas several ED+AD protocols can still achieve finite key rates.
Rather surprisingly, in this noise regime, the protocol $2$-$3$-$1$ performs better than protocol $3$-$2$-$1$. This implies that in the case of depolarizing noise and using a quantum computer that can process at most $3$ qubits as input, using a smaller entanglement distillation ($m=2$) more often ($n=3$) is more advantageous than using a larger entanglement distillation protocol ($m=3$) and performing advantage distillation fewer times ($n=2$). The reason for the rather counter-intuitive situation is as follows: unlike the achievable fidelity calculation, when calculating the achievable key rates, the success probability of the distillation is also important. Although the $3$-$2$-$1$ protocols can, in principle, distill a more entangled state than $2$-$3$-$1$ protocols, the lower overall success probability makes the former less advantageous for performing QKD. In fact, this poses good news for experimental implementation since the larger entanglement distillation protocols require more qubits and longer coherence times in the quantum computer.
@@ IV.2.2 BB84 key rates
In terms of achievable key rates, the ED-only protocol $2$-$1$ outperforms all other more complicated protocols (Fig. 7) in all noise regimes. The special structure of the dephasing noise eliminates the need for complicated distillation protocols such as the $3$-$3$-$1$ protocol. While these protocols can produce high-quality entanglement (Fig. 6), the low success probability makes them less efficient compared to ED-only protocols. The results here demonstrate that depending on the nature of the noise in the input state as well as the strength of the noise, very different distillation protocols should be used in order to achieve optimal performance of entanglement-based QKD.
IV.3 Security bounds for ED+AD distillation protocols
Security bounds are usually defined as the critical quantum bit error rate (QBER) so that for any larger bit error rate, key sharing becomes impossible. It is an important parameter in the security proof of QKD protocols [28]. We first briefly review the previously derived critical QBER for two different QKD protocols. Shor and Preskill proved the unconditional security of BB84 with QBER $<0.11$ with one-way information reconciliation and privacy amplification [18]. It was later improved to $0.124$ by Kraus, Renner, and Gisin with information-theoretical consideration [29, 30]. On the other hand, using advantage distillation is known to improve the security bounds dramatically [13]. The best security bound for BB84 with advantage distillation is $0.2$ [31, 14].
As for the six-state protocol [17] (a variation of the BB84 protocol that uses three mutually unbiased bases instead of two), the best-known security bound with one-way distillation is $0.141$ [30], and the security bound with advantage distillation is $0.276$ [31, 14].
Note that using entanglement distillation before the entanglement-based QKD protocol makes the security condition identical to the entanglement condition since any bi-partite entangled state can always be transformed arbitrarily close to a maximally entangled state given a large enough ED protocol. However, this might be impractical in practice, since entanglement distillation is considered costly.
Here, we derive the security bounds given a small-size ED protocol (for simplicity, we only consider using the DEJMPS protocol) and advantage distillation. In the following analysis, we closely follow the security proof used by Bae and Acin [14]. We first define the Bell-diagonal states $\rho_{\text{Bell}}$ with components $p_{I},p_{X},p_{Y},p_{Z}$ as follow.
$\rho_{\text{Bell}}$ $=p_{I}|\Phi^{+}\rangle\langle\Phi^{+}|$
$+p_{X}|\Psi^{+}\rangle\langle\Psi^{+}|$
$+p_{Y}|\Psi^{-}\rangle\langle\Psi^{-}|$
$+p_{Z}|\Phi^{-}\rangle\langle\Phi^{-}|.$ (4)
In order to share keys, the information-theoretic quantity $I(A:B)-I(A:E)$ must be positive (known as the Csiszár–Körner bound [32]), where $I(A:B)$ is the mutual information of the probability distribution shared by Alice and Bob after a measurement and $I(A:E)$ is the mutual information between Alice and the eavesdropper, Eve. In Ref. [14], the sufficient and necessary condition for sharing keys is derived with the state $\rho_{\text{Bell}}$ shared by Alice and Bob. Given a large but finite round of advantage distillation, the condition is
$\left(p_{I}+p_{Z}\right)\left(p_{X}+p_{Y}\right)<\left(p_{I}-p_{Z}\right)^{2}.$ (5)
Now, we discuss the security bounds for two QKD protocols in detail separately.
IV.3.1 The six-state protocol
We first discuss the critical QBER for the six-state protocol as it is easier to analyze due to its symmetric nature. Similar to the assumption used in Ref. [14], given an observed QBER $Q$, Alice and Bob assume the entangled state $\rho_{AB}$ shared by them is the Werner state with the following components,
$p_{I}=1-\frac{3}{2}Q,$
$p_{X}=\frac{Q}{2},$
$p_{Y}=\frac{Q}{2},$
$p_{Z}=\frac{Q}{2}.$ (6)
Using the state above and evaluating the condition in Eqn. 5 leads to $Q<0.276$, which is the AD-only security bound for the six-state protocol. However, if we apply the DEJMPS protocol to the state in Eqn. IV.3.1 (see appendix B), we obtain the following components,
$\tilde{p}_{I}=\frac{5Q^{2}-6Q+2}{4Q^{2}-4Q+2},$
$\tilde{p}_{X}=\frac{Q^{2}}{4Q^{2}-4Q+2},$
$\tilde{p}_{Y}=\frac{Q^{2}}{4Q^{2}-4Q+2},$
$\tilde{p}_{Z}=\frac{2Q-3Q^{2}}{4Q^{2}-4Q+2}.$ (7)
Allowing permuting the Pauli components (which can be done with local operations only) and evaluating the condition (Eqn. 5) in terms of $Q$ leads to the security bound $Q<0.30$, which outperforms the AD-only security bound.
IV.3.2 The BB84 protocol
Unlike the case of the six-state protocol, in the standard BB84 protocol, only the $z$ and $x$ bases are measured to estimate the error and obtain the raw key. Therefore, one can only assume identical QBER in $x$-basis and $z$-basis but not in $y$-basis. Similar to the analysis done in [14], the state Alice and Bob share is assumed to be as follows.
$p_{I}=1-2Q+x,$
$p_{X}=Q-x,$
$p_{Y}=x,$
$p_{Z}=Q-x,$ (8)
where $x\in[0,Q]$. As Ref. [14] discussed, if Alice and Bob use advantage distillation for the BB84 protocol, the optimal attack by Eve with QBER $Q$ leads to $x=0$. Evaluating the condition in Eqn. 5 leads to $Q<0.2$, which is the AD-only security bound.
However, if Alice and Bob are allowed to perform coherent operations on both sides, they could permute the above components to obtain:
$\tilde{p}_{I}=1-2Q+x,$
$\tilde{p}_{X}=Q-x,$
$\tilde{p}_{Y}=Q-x,$
$\tilde{p}_{Z}=x.$ (9)
Evaluating the security condition Eqn. 5 leads to $Q<0.25$, which is also the entanglement limit QBER for the BB84 protocol [14].
We note that the quantum operations performed by Alice and Bob in the first stage simply permute the coefficients without explicitly performing an entanglement distillation protocol. However, the critical QBER after the permutation of the coefficients is improved to the entanglement limit, which overcomes the AD-only limit ($Q=0.2$). This can be explained by the fact that, in the case of asymmetric noise, the action of permuting Pauli components before the advantage distillation can be viewed as an entanglement distillation protocol [33].
IV.3.3 Mutual information difference with finite rounds of ED and AD
In the previous discussion, we provided the security bounds (as critical QBER) for the six-state protocol and the BB84 protocol using our ED+AD scheme. However, it only tells us the critical QBER such that secret-key generation is possible with a large but finite number of advantage distillation rounds. In practice, Alice and Bob would refrain from using many rounds of advantage distillation as they decrease the overall key rates significantly. Therefore, we study the potentially achievable key rates with finite rounds of ED and AD. From this information, we can learn the minimal size of the distillation protocols in order to reach a positive key rate under a given noise level.
Specifically, we calculate the mutual information difference $I(A:B)-I(A:E)$ with finite rounds of entanglement distillation and advantage distillation. This has been derived previously in Ref. [14].
$I(A:B)-I(A:E)=$ $1-H(\epsilon_{N})-(1-\epsilon_{N})H\left(\frac{1-\Lambda^{N}_{\text{eq}}}{2}\right)$
$-\epsilon_{N}H\left(\frac{1-\Lambda^{N}_{\text{diff}}}{2}\right),$ (10)
where $N$ is the number of AD rounds, and
$\epsilon_{N}$ $=\frac{\epsilon^{N}_{AB}}{\epsilon^{N}_{AB}+(1-\epsilon_{AB})^{N}},$
$\epsilon_{AB}$ $=p_{X}+p_{Y},$
$\Lambda_{\text{eq}}$ $=\frac{p_{I}-p_{Z}}{p_{I}+p_{Z}},$
$\Lambda_{\text{diff}}$ $=\frac{\lvert p_{X}-p_{Y}\rvert}{p_{X}+p_{Y}}.$ (11)
We summarize the results in Fig. 8. For both BB84 and the six-state protocol, we calculate the mutual information difference (Eqn. 10) for using different finite combinations of ED and AD. For each choice of entanglement distillation protocol, we plot the maximum and minimum of the quantities $I(A:B)-I(A:E)$ considering up to $10$ rounds of advantage distillation (solid lines). The shaded area between the solid lines is the achievable mutual information difference with a specific ED protocol given a fixed QBER.
For both BB84 and the six-state protocol, we consider three cases of using entanglement distillation (described in the caption). The non-overlapping area with increasing number of entanglement distillation indicates the advantage of using more a powerful entanglement distillation protocol with the advantage distillation. Most importantly, the results imply using a modest number of distillation resources (EDx2 for six-state protocol and permuting the coefficients for BB84 with AD up to 10 rounds) can already lead to critical QBER greater than the previously derived critical QBER ($0.2$ for BB84 and $0.276$ for the six-state protocol). In practice, Alice and Bob could estimate the QBER in the parameter estimation scheme of the protocol. Based on the QBER information, they decide the optimal ED+AD protocols in order to distill the secret keys.
The results here demonstrate the interplay between entanglement distillation, advantage distillation, and the key rate for a specific QKD protocol with our ED+AD distillation scheme and provide an estimation of the resources required for the key distillation.
Figure 8: Achievable key rates with finite-size ED+AD distillation. (left) For the six-state protocol, we consider the cases of not using entanglement distillation (EDx0), using the DEJMPS protocol once (EDx1) and using the DEJMPS protocol twice (EDx2). (right) For the BB84 protocol, we consider the cases of not using entanglement distillation (EDx0), permuting the Pauli coefficients (EDx0 (permute)), and using the DEJMPS protocol once (EDx1). For both protocols, we calculate the achievable mutual information difference within 10 rounds of advantage distillation and plot as the shaded area.
@@ V Conclusion
We presented a scheme for improving entanglement-based quantum key distribution based on entanglement distillation and advantage distillation. We find both increased secret-key rates and increased resilience against noise while not imposing any additional requirements on the quantum resources. In particular, our protocol has the potential to achieve a non-zero key rate even at a noise regime where using entanglement distillation only leads to a vanishing key rate for the case of depolarizing noise.
We derived security bounds for the ED+AD distillation scheme which outperforms the previous security bounds using AD-only key distillation protocols. Moreover, the improved tolerance to the noise can be achieved within modest-size distillation protocols.
Our approach also works for more general Pauli noise models — if one possesses knowledge about the initial noisy state shared by two parties, our framework can be immediately utilized to find the optimal ED+AD protocol for performing entanglement-based QKD. For example, applying our tools to the case of pure-dephasing noise, we found that the basic $2$-$1$ protocol achieves better key rates than more sophisticated protocols, further highlighting the practical implications of our results on near-term quantum communication.
Our work suggests several possible future directions of study. First, due to the still rather large number of all possible ED+AD protocols, studying the $m$-$n$-$1$ protocol for large $m$ and $n$ is still formidable. A discrete optimization algorithm, rather than enumerating all possible protocols, should be employed to study larger protocols. Second, we have not considered the case of more general $m$-$n$-$o$ ED+AD protocols ($n>o>1$). In this more general case, one would need to consider all possible $n$-$o$ advantage distillation protocols to obtain optimal protocols. Finally, in practice, one has to deal with non-identical input states and the effect of imperfect quantum operations during entanglement distillation [16]. We believe the framework built here can be extended to take the above considerations in mind, which in turn will yield more efficient entanglement-based QKD protocols and applications in the near future.
@@ Appendix A Binary picture and the symplectic representation of Clifford operators
For enumerating the bi-local Clifford protocols, it will be convenient to define an equivalence relation on such protocols; equivalent protocols have equivalent distillation statistics, meaning the success probability and output state (up to local rotations) are the same. Here, we review the required mathematical framework. The material here follows Ref. [15] closely.
In the study of Clifford distillation protocols, the following binary vector representation for the Pauli group will be useful. The convention here follows Ref. [34] and Ref. [15]. We define the following mapping between binary vectors and single-qubit Pauli operators as follows
@@
Nevertheless, when one considers using Clifford gates to perform entanglement distillation, not all choices of Clifford operators lead to different distillation statistics. Two protocols have the same distillation statistics if the output fidelity $F=\langle\Phi^{+}|\rho_{distilled}|\Phi^{+}\rangle$ and the success probability are the same, and other components with respect to the Bell basis ($p_{X}=\langle\Psi^{+}|\rho_{\text{distilled}}|\Psi^{+}\rangle,p_{Y}=\langle\Psi^{-}|\rho_{\text{distilled}}|\Psi^{-}\rangle,p_{Z}=\langle\Phi^{-}|\rho_{\text{distilled}}|\Phi^{-}\rangle$) are the same up to some permutation. In Ref. [15], the equivalence classes (up to the same distillation statistics) of the symplectic group are characterized, which has a much smaller size than the original symplectic group. Let ${D}_{n}$ be an element in the symplectic group $\text{Sp}(2n,\mathbb{Z}_{2})$ which, when left multiplied to any Clifford operator $C_{n}$ in the distillation protocol, does not change the distillation statistics. Such ${D}_{n}$ form a subgroup $\mathcal{D}_{n}$. When one wants to enumerate all possible bi-local Clifford protocols for a given input state, it suffices to enumerate one element in each coset of $\mathcal{D}_{n}$ in $\text{Sp}(2n,\mathbb{Z}_{2})$, which has a size equal to the associated subgroup index
@@
Appendix B The DEJMPS protocol with Bell diagonal input states
Here, we explicitly show the output state of the DEJMPS entanglement distillation protocol (see Sec. II.2 given an input in the Bell diagonal form. (Eqn. IV.3)
We assume the input states are two identical Bell states. Hence, the state before the entanglement distillation is $\rho^{\otimes 2}_{\text{Bell}}$, with components $\rho^{\otimes 2}_{\text{Bell},ij}=\left(\rho_{\text{Bell,i}}\right)\left(\rho_{\text{Bell,j}}\right)$ for $i,j\in{I,X,Y,Z}$. Upon success, the components of the output state $\tilde{\rho}$ and the successful probability $p_{\text{suc}}$ are
$p_{\text{suc}}$ $=p_{II}+p_{XX}+p_{ZX}+p_{YI}$
$+p_{YY}+p_{ZZ}+p_{XZ}+p_{IY},$
$\tilde{\rho}_{I}$ $=\frac{\left(p_{II}+p_{YY}\right)}{p_{\text{suc}}},$
$\tilde{\rho}_{X}$ $=\frac{\left(p_{XX}+p_{ZZ}\right)}{p_{\text{suc}}},$
$\tilde{\rho}_{Y}$ $=\frac{\left(p_{ZX}+p_{XZ}\right)}{p_{\text{suc}}},$
$\tilde{\rho}_{Z}$ $=\frac{\left(p_{YI}+p_{IY}\right)}{p_{\text{suc}}}.$ (17)
Note that in order to maximize the output fidelity $\tilde{\rho}_{I}$, one should permute the input state’s components such that $p_{I}>p_{X}\geq p_{Z}\geq p_{Y}$ [36].
Appendix C Comparison with DEJMPS-based protocols
@@
Figure 9: Illustration of a $3$-$1$ DEJMPS protocol. The $3$-$1$ DEJMPS protocol is the recursive version of the standard DEJMPS protocol. The first unmeasured pair is the output of the protocol.
@@
Appendix D Repetition code based distillation on Werner states
@@ Acknowledgements.
We thank Hiroki Takahashi for stimulating discussions. S. S. thanks Xiongfeng Ma, William J. Munro, and Koji Azuma for their comments regarding the security bounds. This project was supported by the JST Moonshot R&D program under Grant JPMJMS226C.